# Mitigating photon loss in linear optical quantum circuits: classical postprocessing methods outperforming postselection

### v2

@@ 5.2 Decomposition of recycled probabilities into $n$–photon output bit string probabilities
where the sum includes all the bit string outputs $s^{n}_{j}\in\mathcal{G}(s_{i}^{n-k})$ from which the loss of $k$ photons maps to $s_{i}^{n-k}$. Each element of the sum is composed of the probability $p(s^{n}_{j})$ that the output is the $n$–photon bit string $s^{n}_{j}$, multiplied by the uniform probability (${{n\choose k}}^{-1}$ of the loss of $k$ photons from $s^{n}_{j}$ resulting in the output $s^{n-k}_{i}$.
As in eqn. 6, in the definition of the recycled probabilities in eqn. 4 each of the lossy probabilities can be decomposed into a convex combination of probabilities from the ideal distribution. That is, the $n-k$–photon output probabilities within the sum in eqn. 4 can be replaced with a sum over $n$–photon output probabilities. For some arbitrary labelling of bit strings $s^{n}_{j}$, and noting that $s^{n}_{l}\in\mathcal{G}(s_{i}^{n-k})$, we can expand the $n-k$ output photon probabilities in the form
@@
$\begin{split}p_{R}^{k}(s^{n}_{l})&=p(s^{n}_{l})+\sum_{s_{i}^{n-k}\in\mathcal{L}(s^{n}_{l})}\sum_{s^{n}_{j}\in\mathcal{G}(s_{i}^{n-k}),j\neq l}p(s^{n}_{j})\frac{1}{{n\choose k}}.\\ \end{split}$ (8)
@@
$\begin{split}\sum_{l=1}^{{m\choose n}}p_{R}^{k}(s^{n}_{l})&=\sum_{l=1}^{{m\choose n}}\sum_{s_{i}^{n-k}\in\mathcal{L}(s^{n}_{l})}p(s^{n-k}_{i})\\ &=\sum_{l=1}^{{m\choose n}}\sum_{s_{i}^{n-k}\in\mathcal{L}(s^{n}_{l})}\sum_{s^{n}_{j}\in\mathcal{G}(s_{i}^{n-k})}p(s^{n}_{j})\frac{1}{{n\choose k}}\\ &={{m-n+k}\choose k},\\ \end{split}$ (9)
This is because $\sum_{l}p(s^{n}_{l})=1$, and, because $|\mathcal{L}(s^{n}_{i})|={n\choose k}$ and $|\mathcal{G}(s^{n-k}_{i})|={m-n+k\choose k}$, each distinct bit string $s^{n}_{l}$ appears exactly (${n\choose k}{m-n+k\choose k}$ times in the above sum  [^2] Another way to see this is to consider the sum over recycled probabilities as $\sum_{l}p_{R}(s^{n}_{l})=\sum_{l}\sum_{s_{i}^{n-k}\in\mathcal{L}(s^{n}_{l})}p(s^{n-k}_{i})={m-n+k\choose k}$. Where the rightmost equality follows from the fact that each $p(s^{n-k}_{i})$ appears exactly (${m-n+k\choose k}$ times in this sum, since there are $|\mathcal{G}(s^{n-k}_{i})|$ recycled probabilities sharing a single $p(s^{n-k}_{i})$. Furthermore note that $\sum_{i}p(s^{n-k}_{i})=1$. . Therefore $\mathbf{N}=\frac{1}{{m-n+k\choose k}}$, and the expression for the normalised recycled probability is
$\begin{split}p_{R}^{k}(s^{n}_{l})&=p(s^{n}_{l})\frac{1}{{m-n+k\choose k}}+\sum_{s_{i}^{n-k}\in\mathcal{L}(s^{n}_{l})}\sum_{s^{n}_{j}\in\mathcal{G}(s_{i}^{n-k}),j\neq l}p(s^{n}_{j})\frac{1}{{m-n+k\choose k}{n\choose k}}.\\ \end{split}$ (10)
@@
Importantly, the photon recycled probability $p_{R}^{k}(s^{n}_{l})$ contains an amplified signal of the ideal probability $p(s^{n}_{l})$, relative to the other probabilities contained within the interference term.
6 Properties of the recycled distribution
As mentioned in the introduction section, recycling mitigation performs best in the regime of high loss $\eta>0.5$, and when using $n-k$ output photon statistics where $k$ is low to construct the recycled probabilities. In this section, we show that recycled probabilities constructed from output statistics where most photons were lost (i.e. when $n-k$ is a constant independent of $n$) are efficiently classically computable, and therefore not useful for obtaining interesting mitigation performance. Furthermore, we provide evidence that recycled probabilities constructed from $n-k$ output statistics where $k$ is a constant independent of $n$ (corresponding to output statistics where a small number of photons have been lost) are hard to compute classically, making these interesting to use for recycling mitigation protocols.
@@ 6 Properties of the recycled distribution
where the number of terms of this sum is (${n\choose k}$. For Haar-random interferometers $U$, in the no-collision regime, we can use the results of [17], and in turn express each $p(s^{n-k}_{i})$ as
@@
where the $X_{i}$’s are $n-k\times n-k$ matrices with independently distributed Gaussian entries [6], and the number of terms in this sum is also (${n\choose k}$. We show that in the high loss regime, where $k=n-r$, and $r$ is a constant independent of $n$, the sum $\sum_{s_{i}^{n-k}\in\mathcal{L}(s^{n}_{l})}p(s^{n-k}_{i})$ is computable efficiently classically. Our result is encompassed in the following lemma proven in appendix 15.
@@
Let $k=n-r$, there is a classical algorithm running in time $O(2^{r-1}r\big{(}{n\choose n-r}\big{)}^{2})$ which exactly computes $\sum_{s_{i}^{n-k}\in\mathcal{L}(s^{n}_{l})}p(s^{n-k}_{i})$.
Notice that when $r$ is a constant independent of $n$, the runtime of the above algorithm is $O({n\choose n-r})\leq O(n^{r})\leq\mathsf{poly}(n)$. Thus, recycled probabilities corresponding to a high number of lost photons are efficiently computable classically. For low values of loss, in particular when $r$ scales with $n$, the above efficient classical simulability results break down, this however does not necessarily imply that there is no other, possibly efficient, algorithm for computing $\sum_{s_{i}^{n-k}\in\mathcal{L}(s^{n}_{l})}p(s^{n-k}_{i})$.
For the case where $k$ is a constant independent of $n$, there is evidence that computing $\sum_{s_{i}^{n-k}\in\mathcal{L}(s^{n}_{l})}p(s^{n-k}_{i})$ is probably hard (inefficient) to do classically. Indeed, it is known that in worst-case the probabilities $p(s^{n-k}_{i})$ when $k$ is a constant independent of $n$ are not efficient to compute classically, unless the polynomial hierarchy collapses to its third level [17]. This motivates the application of recycling mitigation in the low $k$ regime, as the $n-k$–photon output statistics are efficiently computable classically when $k$ is high, and therefore any potential quantum advantage is lost. In all our simulations, we apply our error mitigation techniques to statistics where the number of lost photons is a constant independent of system size, and discard all other statistics.
7 Bounding deviation of interference terms
Methods are later presented to extract the ideal probability signals from the recycled probabilities. The precision with which this operation may be achieved is limited by variation of the interference terms from their expected value. As such, upper bounding this variation behaviour is crucial when we later assess the accuracy of the mitigated values. The recycled probabilities have a composite structure, comprising a mixture of the ideal probability and the interference term. While the interference term,
@@ 7 Bounding deviation of interference terms
is itself a mixture of $n$–photon output probabilities. It is possible to account for the errors caused by using approximations of the interference terms when generating the mitigated probabilities by upper bounding the deviation of the interference term away from an expected value.
@@
$\begin{split}\mathbf{E}_{X\in\mathcal{G}_{n\times n}}\big{(}I_{s^{n}_{l},k}\big{)}&=p_{unif}\big{(}1+O(m^{-1})\big{)}\approx p_{unif}.\\ \end{split}$ (14)
with $\mathbf{E}_{X\in\mathcal{G}_{n\times n}}(.)$ the expectation value over the set $\mathcal{G}_{n\times n}$.
@@
$\begin{split}\text{Pr}\Big{(}\big{|}I_{s^{n}_{l},k}-p_{unif}\big{|}\geq\epsilon_{bias,s^{n}_{l}}\Big{)}&\leq\frac{p_{unif}}{\epsilon_{bias,s^{n}_{l}}^{2}},\\ \end{split}$ (17)
@@
While both these upper bounds use Chebyshev’s inequality [27], the Thm. 3 bound applies only to Haar random matrices, while the Thm. 5 bound applies to arbitrary matrices. However, to upper bound the variance for the Haar random case we use exact moments originally derived in [61]. While in the case of arbitrary matrices the Bhatia-Davis inequality [62] is instead used, which results in a looser bound.
We note that a tighter, although distribution-dependent, bound may be found than the one provided in Thm. 5. This result follows from Lemmas 6 and 7, proven in appendix 17, which will now be stated, that use the definition of the variance that for a set of real values $\{x_{i}\}_{i}$ with mean $\mu$ and cardinality $N$, $\mathsf{Var}(\{x_{i}\}_{i}):=N^{-1}\sum_{i}(x_{i}-\mu)^{2}$.
@@
Using these lemmas, an upper bound on the largest probability of the output distribution may be used to derive a tighter upper bound on the variance using the Bhatia-Davis inequality [62]. Let $p_{\text{upper}}$ be an experimentally derived upper bound on the largest probability in the ideal $n$–photon output distribution $p_{\text{max}}$, such that $p_{\text{upper}}\geq p_{\text{max}}$. Following similar steps as for the proof for Thm. 5 results in an upper bound on the confidence of $1-\frac{p_{unif}p_{{upper}}}{\epsilon_{bias,s^{n}_{l}}^{2}}-\delta\Big{(}1-\frac{p_{unif}p_{{upper}}}{\epsilon_{bias,s^{n}_{l}}^{2}}\Big{)}$, where $\delta$ is the confidence parameter for the $p_{upper}$ estimator. This result is stated formally in the following.
@@
$\begin{split}\text{Pr}\Big{(}\big{|}I_{s^{n}_{l},k}-p_{unif}\big{|}\geq\epsilon_{bias,s^{n}_{l}}\Big{)}&\leq\frac{p_{unif}p_{{upper}}}{\epsilon_{bias,s^{n}_{l}}^{2}}+\delta\Big{(}1-\frac{p_{unif}p_{{upper}}}{\epsilon_{bias,s^{n}_{l}}^{2}}\Big{)},\\ \end{split}$
@@
A proof of this result is given in appendix 17. We note also that as $p_{\text{max}}\leq 1$ with confidence $1$, so that $\delta=0$ in Thm. 8, Thm. 5 follows as a corollary.
With assumptions about the structure of the unitaries, like in [28], it is possible to deterministically and exponentially upper bound the interference terms. For a large class of unitary matrices, we show that the bias error scales as an inverse exponential in $n$. As will be seen later, this shows that our mitigation techniques outperform postselection in estimating output probabilities and expectation values of linear optical circuits for up to inverse exponential precisions. Here we use the the operator 2-norm and the infinity norm. For an $n\times n$ matrix $A$ the operator 2-norm is defined $\|A\|_{2}:=\sup_{\|\vec{x}\|_{2}\leq 1,\vec{x}\in\mathbb{C}^{n}}\|A\vec{x}\|_{2}$, where $\|\vec{v}\|_{p}$ is the $l_{p}$ norm, that is $\|\vec{v}\|_{p}=\big{(}\sum_{p}|v_{i}|^{p}\big{)}^{1/p}$. And the infinity norm which is defined as $\|\vec{v}\|_{\infty}:=\max_{i}|v_{i}|$. Let $h^{A}_{\infty}:=\frac{1}{n}\sum_{i=1,\dots,n}\|\mathbf{A}_{i}\|_{\infty}$, where $\mathbf{A}_{i}$ is the $i$th row of $A$. In appendix 20 we show the following.
@@
$\epsilon_{bias,s^{n}_{l}}\leq\approx O(e^{-0.000002n}).$ (20)
In other words, for the stated class of unitary matrices the bias error is inverse-exponentially (in $n$) close to zero. For this class of unitary matrices, one can show, as will be seen later on, that our mitigation techniques outperform postselection for up to inverse-exponential additive errors. However the number of photons required for this bound to become significant ($n\approx 10^{5}$) is beyond what is possible with current technology.
We will now apply the derived bounds to give performance guarantees for several methods which may be applied to extract the signal of the ideal probabilities from the recycled probabilities. In the following, to make the bounds in Thms. 3 and 5 more concrete we will set $\epsilon_{bias,s^{n}_{l}}=O\Big{(}\frac{1}{\mathsf{poly}(m)}\Big{)}$.
8 Generating the loss-mitigated outputs
We now present two methods for constructing loss mitigated outputs from the recycled probabilities, these we refer to as linear solving and extrapolation. In linear solving, the interference term within each recycled probability is substituted for its expected value, and the resulting expressions are then solved to find estimators of the ideal probabilities. While in extrapolation, the decay of the ideal signal in the set of recycled distributions with $k$ is used to compute estimators of the ideal probabilities. We derive inequalities that are conditions for recycling mitigation with the relevant postprocessing techniques to beat postselection.
The output photon statistics where most photons have been lost are not used in recycling mitigation. Indeed, it is clear that if one uses output photon statistics with high $k$ to construct the recycled probabilities, then these statistics may be efficiently classically simulated; as per Lemma 1. This strongly motivates the construction of recycled distributions from output statistics where $k$ is low.
@@ 8 Generating the loss-mitigated outputs
As the bias errors in the bounds for Thm. 3 and Thm. 5 have been set to $\epsilon_{bias,s^{n}_{l}}=O\big{(}\frac{1}{\mathsf{poly}(m)}\big{)}$, the performance guarantees for the Haar random and arbitrary matrix cases are of the same form. Therefore, for the sake of concision we state one performance guarantee for each version of the classical postprocessing. Note, however, that the confidence for the Haar random matrix bias error upper bound is $\approx 1-O^{*}(p_{unif}^{-2})$, whereas for the arbitrary matrix bias error upper bound is $\approx 1-O^{*}(p_{unif}^{-1})$  [^3] The $O^{*}(f)$ notation means a complexity of $O(f)$ up to polynomial prefactors, that is $O^{*}(f):=O(\mathsf{poly}(n)f)$. Note that, since $m=O(\mathsf{poly}(n))$ for any useful application, then $O^{*}(f)$ can equivalently be thought of as $O(f)$ up to a $\mathsf{poly}(m)$ prefactor. .
8.1 Linear solving
The linear solving method involves substituting the interference term within each recycled probability for an approximate value, and then solving the resulting expressions for the ideal probabilities. As the expectation of interference terms over the Haar measure and over the recycled distribution are $p_{unif}+O(m^{-n+1})$ and $p_{unif}$, respectively, see Lemmas 2 and 4, the term $\frac{N^{\prime}_{k}}{N_{k}}p_{unif}$ is used for this substitution. This allows the upper bounding of the error introduced by the substitution using Thms. 3 and 5. Each recycled probability constructed from the $n-k$ output photon statistics may then be written

### v3

@@ 4.1.2 Decomposition of recycled probabilities into nn–photon output bit string probabilities
where the sum includes all the bit string outputs $s^{n}_{j}\in\mathcal{G}(s_{i}^{n-k})$ from which the loss of $k$ photons maps to $s_{i}^{n-k}$. Each element of the sum is composed of the probability $p(s^{n}_{j})$ that the output is the nn–photon bit string $s^{n}_{j}$, multiplied by the uniform probability ${{n\choose k}}^{-1}$ of the loss of $k$ photons from $s^{n}_{j}$ resulting in the output $s^{n-k}_{i}$.
As in eqn. 12, in the definition of the recycled probabilities in eqn. 10 each of the lossy probabilities can be decomposed into a convex combination of probabilities from the ideal distribution. That is, the nkn-k–photon output probabilities within the sum in eqn. 10 can be replaced with a sum over nn–photon output probabilities. For some arbitrary labelling of bit strings $s^{n}_{j}$, and noting that $s^{n}_{l}\in\mathcal{G}(s_{i}^{n-k})$, we can expand the nkn-k output photon probabilities in the form
@@
[…]
@@
[…]
This is because $\sum_{l}p(s^{n}_{l})=1$, and, because $|\mathcal{L}(s^{n}_{i})|={n\choose k}$ and $|\mathcal{G}(s^{n-k}_{i})|={m-n+k\choose k}$, each distinct bit string $s^{n}_{l}$ appears exactly ${n\choose k}{m-n+k\choose k}$ times in the above sum  [^2] Another way to see this is to consider the sum over recycled probabilities as $\sum_{l}p_{R}(s^{n}_{l})=\sum_{l}\sum_{s_{i}^{n-k}\in\mathcal{L}(s^{n}_{l})}p(s^{n-k}_{i})={m-n+k\choose k}$. Where the rightmost equality follows from the fact that each $p(s^{n-k}_{i})$ appears exactly $m-n+k\choose k$} times in this sum, since there are $|\mathcal{G}(s^{n-k}_{i})|$ recycled probabilities sharing a single $p(s^{n-k}_{i})$. Furthermore note that $\sum_{i}p(s^{n-k}_{i})=1$. . Therefore $\mathbf{N}=\frac{1}{{m-n+k\choose k}}$, and the expression for the normalised recycled probability is
[…]
@@
Importantly, the photon recycled probability $p_{R}^{k}(s^{n}_{l})$ contains an amplified signal of the ideal probability $p(s^{n}_{l})$, by a factor of $n\choose k$}, relative to the other probabilities contained within the interference term.
4.2 A classical simulation algorithm for recycled probabilities
In this section, we show that recycled probabilities constructed from output statistics where most photons were lost (i.e. when nkn-k is a constant independent of $n$) are efficiently classically computable, and therefore not useful for obtaining interesting mitigation performance. Furthermore, we provide evidence that recycled probabilities constructed from nkn-k output statistics where $k$ is a constant independent of $n$ (corresponding to output statistics where a small number of photons have been lost) are hard to compute classically, making these interesting to use for recycling mitigation protocols.
@@ 4.2 A classical simulation algorithm for recycled probabilities
where the number of terms of this sum is $n\choose k$}. For Haar-random interferometers $U$, in the no-collision regime where $m=\Omega(n^{5})$, we can use the results of [3], and in turn express each $p(s^{n-k}_{i})$ as
@@
where the $X_{i}$’s are $n-k\times n-k$ matrices with independently distributed Gaussian entries [1], and the number of terms in this sum is also $n\choose k$}. We show that in the high loss regime, where k=nrk=n-r, and $r$ is a constant independent of $n$, the sum $\sum_{s_{i}^{n-k}\in\mathcal{L}(s^{n}_{l})}p(s^{n-k}_{i})$ is computable efficiently classically. Our result is encompassed in the following lemma proven in appendix 9.
@@
Let k=nrk=n-r, there is a classical algorithm running in time $O(2^{r-1}r\big({n\choose n-r}\big)^{2})$ which exactly computes $\sum_{s_{i}^{n-k}\in\mathcal{L}(s^{n}_{l})}p(s^{n-k}_{i})$.
Notice that when $r$ is a constant independent of $n$, the runtime of the above algorithm of the order of ${n\choose n-r}\in O(n^{r})$ which is a polynomial in $n$. Thus, recycled probabilities corresponding to a high number of lost photons are efficiently computable classically. For low values of loss, in particular when $r$ scales with $n$, the above efficient classical simulability results break down, this however does not necessarily imply that there is no other, possibly efficient, algorithm for computing $\sum_{s_{i}^{n-k}\in\mathcal{L}(s^{n}_{l})}p(s^{n-k}_{i})$.
For the case where $k$ is a constant independent of $n$, there is evidence that computing $\sum_{s_{i}^{n-k}\in\mathcal{L}(s^{n}_{l})}p(s^{n-k}_{i})$ is probably hard (inefficient) to do classically (see also the discussion in Section 6). Indeed, it is known that in worst-case the probabilities $p(s^{n-k}_{i})$ when $k$ is a constant independent of $n$ are not efficient to compute classically, unless the polynomial hierarchy collapses to its third level [3]. This motivates the application of recycling mitigation in the low $k$ regime, as the nkn-k–photon output statistics are efficiently computable classically when $k$ is high, and therefore any potential quantum advantage is lost. In all our simulations, we apply our error mitigation techniques to statistics where the number of lost photons is a constant independent of system size, and discard all other statistics.
4.3 Bounding the bias error
In this section, we give bounds on the bias error induced when replacing the interference term with its expectation value. This bias error will be present in all the error mitigation techniques we derive later on. We provide probabilistic bounds on this error, using statistical inequalities. Also, we provide a deterministic upper bound by using the results of [8] which hold for specific families of unitary matrices.
The recycled probabilities have a composite structure, comprising a mixture of the ideal probability and the interference term. While the interference term,
@@ 4.3 Bounding the bias error
is itself a mixture of nn–photon output probabilities. It is possible to account for the errors caused by using approximations of the interference terms when generating the mitigated probabilities by upper bounding the deviation of the interference term away from an expected value.
@@
[…]
where $g(m)\in O(m^{-1})$, with $\mathbf{E}_{X\in\mathcal{G}_{n\times n}}(.)$ the expectation value over the set $\mathcal{G}_{n\times n}$.
@@
[…]
@@
While both these upper bounds use Chebyshev’s inequality [37], the Thm. 8 bound applies only to Haar random matrices, while the Thm. 10 bound applies to arbitrary matrices. However, to upper bound the variance for the Haar random case we use exact moments originally derived in [46]. While in the case of arbitrary matrices the Bhatia-Davis inequality [9] is instead used, which results in a looser bound.
We note that a tighter, although distribution-dependent, bound may be found than the one provided in Thm. 10. This result follows from Lemmas 11 and 12, proven in Appendix 11, which will now be stated, that use the definition of the variance that for a set of real values $\{x_{i}\}_{i}$ with mean $\mu$ and cardinality $N$, $\mathsf{Var}(\{x_{i}\}_{i}):=N^{-1}\sum_{i}(x_{i}-\mu)^{2}$.
@@
Using these lemmas, an upper bound on the largest probability of the output distribution may be used to derive a tighter upper bound on the variance using the Bhatia-Davis inequality [9]. Let $p_{\text{upper}}$ be an experimentally derived upper bound on the largest probability in the ideal nn–photon output distribution $p_{\text{max}}$, such that $p_{\text{upper}}\geq p_{\text{max}}$. Following similar steps as for the proof for Thm. 10 results in an upper bound on the confidence of $1-\frac{p_{unif}p_{{upper}}}{\epsilon_{bias,s^{n}_{l}}^{2}}-\delta\Big(1-\frac{p_{unif}p_{{upper}}}{\epsilon_{bias,s^{n}_{l}}^{2}}\Big)$, where $\delta$ is the confidence parameter for the $p_{upper}$ estimator. This result is stated formally in the following.
@@
[…]
@@
A proof of this result is given in appendix 11. We note also that as $p_{\text{max}}\leq 1$ with confidence $1$, so that $\delta=0$ in Thm. 13, Thm. 10 follows as a corollary.
With assumptions about the structure of the unitaries, like in [8], it is possible to deterministically and exponentially upper bound the interference terms.
For a large class of unitary matrices, we show that the bias error scales as an inverse exponential in $n$. As will be seen later, this shows that our mitigation techniques outperform postselection in estimating output probabilities and expectation values of linear optical circuits for up to inverse exponential precisions. Here we use the the operator 2-norm and the infinity norm. For an $n\times n$ matrix $A$ the operator 2-norm is defined $\|A\|_{2}:=\sup_{\|\vec{x}\|_{2}\leq 1,\vec{x}\in\mathbb{C}^{n}}\|A\vec{x}\|_{2}$, where $\|\vec{v}\|_{p}$ is the $l_{p}$ norm, that is $\|\vec{v}\|_{p}=\big(\sum_{p}|v_{i}|^{p}\big)^{1/p}$. And the infinity norm which is defined as $\|\vec{v}\|_{\infty}:=\max_{i}|v_{i}|$. Let $h^{A}_{\infty}:=\frac{1}{n}\sum_{i=1,\dots,n}\|\mathbf{A}_{i}\|_{\infty}$, where $\mathbf{A}_{i}$ is the $i$th row of $A$. In Appendix 14, we use a result of [8] to show the following.
@@
$\epsilon_{bias,s^{n}_{l}}\in O(e^{-2\times 10^{-5}n}).$ (26)
Interestingly, the probabilistic bounds of Theorems 8 and 10 also give, with high-confidence, exponentially low bounds on the bias error. Indeed, by setting for example $\epsilon_{bias,s^{n}_{l}}\in\Omega(\sqrt{p_{unif}})$ in Theorem 10, we observe that $|I_{s^{n},k}-p_{unif}|$ is upper bounded by $\Omega(\sqrt{p_{unif}})$ with confidence 1o(1)1-o(1). Intuitively, and as will be seen later in more details, both the probabilistic and deterministic upper bounds on the bias error indicate that recycling mitigation outperforms post-selection for up to inverse-exponential in $n$ precision.
Appendix 16.1 discusses possible directions to improve our performance guarantees, and contains a technical result about sums of permanents of i.i.d. Gaussian matrices [1] that might find use beyond this work.
4.4 Generating the loss-mitigated outputs
We now present two methods for constructing loss mitigated outputs from the recycled probabilities, these we refer to as linear solving and extrapolation. In linear solving, the interference term within each recycled probability is substituted for its expected value, and the resulting expressions are then solved to find estimators of the ideal probabilities. While in extrapolation, the decay of the ideal signal in the set of recycled distributions with $k$ is used to compute estimators of the ideal probabilities.
@@ 4.4 Generating the loss-mitigated outputs
4.4.1 Linear solving
The linear solving method involves substituting the interference term within each recycled probability for an approximate value, and then solving the resulting expressions for the ideal probabilities. As the expectation of interference terms over the Haar measure and over the recycled distribution are $p_{unif}+O(m^{-n+1})$ and $p_{unif}$, respectively, see Lemmas 7 and 9, the term $\frac{N^{\prime}_{k}}{N_{k}}p_{unif}$ is used for this substitution. This allows the upper bounding of the error introduced by the substitution using Thms. 8 and 10. Each recycled probability constructed from the nkn-k output photon statistics may then be written